Which equations have infinitely many solutions?

5x = 5(x-1) + 2


3x - 2x + 4 = 2 + x + 2


11 - 6x = 2(3x + 4)


7x - 2 = 7(x - 2)


12 - 8x = 2(5x + 7)


3x - 4 = 3(x - 4) + 8
Answer #1

ANSWER ;


When x = x , then infinitely many solutions are there.


Accordingly, following equations have infinitely many solutions :


1.  3x - 2x + 4 = 2 + x + 2


2. 3x - 4 = 3(x - 4) + 8 

answered by: Tulsiram Garg